# TEASMA: A Practical Approach for Test Assessment of Deep Neural Networks using Mutation Analysis

### v2

@@ VI-A RQ1: Can an accurate regression model be built to explain FDR as a function of MS?
Figure 4: Linear regression line between $\widehat{FDR}$ and $ActualFDR$
VI-C Threats to Validity
@@
Internal threats to validity One of the internal threats to validity pertains to our sampling approach, as our results are based on sampled input subsets. To address this, we selected diverse input subsets, employing both random and uniform sampling techniques and varying sample sizes. Another internal threat to validity stems from selected parameters used throughout the experiments. This encompasses parameters for clustering mispredicted inputs to estimate faults and for configuring MOs. To mitigate this concern, regarding clustering parameters, we explored different settings for each subject and opted for the best configuration based on the evaluation of the Silhouette index [34] and DBCV scores [35], as suggested by Aghababaeyan et al. [6]. For MO configuration, we used the default settings of the DeepMutation++ tool, which yielded accurate results in terms of FDR prediction.
@@
Construct threats to validity A potential threat is related to the calculation of MS and FDR. To mitigate this concern and ensure the robustness of our results, we considered multiple ways to compute MS, as described in Section II-B. Regarding FDR, we leveraged a SoTA fault estimation approach [6] based on clustering mispredicted inputs. This approach has been thoroughly validated and the resulting clusters have been carefully analyzed by the authors. We reused their publicly available approach to achieve accurate fault identification.
@@
External threats to validity To address potential concerns regarding the generalizability of TEASMA, we performed our experiment on multiple subjects including various combinations of five DNN models with different internal architectures and four input sets containing diverse types of images. Another threat might be related to the specific set of MOs that we considered in our experiment. For practical reasons, we focused on post-training MOs and investigated all the MOs implemented by the DeepMutation++ tool, also considering whether subsets of MOs might yield better results. Other MOs may yield different results in the future.
VII The Cost of Using Pre-training Mutants
In this section, we compare the cost associated with performing mutation analysis based on pre-training mutants generated by DeepCrime with post-training mutants generated by DeepMutation++. Though this is not the primary focus of this paper, it provides initial evidence supporting our rationale for employing post-training MOs in TEASMA and our concerns regarding the impractical cost of pre-training mutants, as previously outlined in Section II-C. To perform this experiment we used the same computing environment described in Section IV-C. Given the experiment’s time-intensive nature, which entails consuming enormous amounts of computing resources, conducting this experiment on all our subjects would take many weeks and is neither feasible nor necessary for our purpose. This problem is made more acute by unpredictable server outages, which are unavoidable in such large experiments. As a result, we opted to focus on one of our smaller DNNs, specifically subject S1 (LeNet5), since it is expected to provide a conservative result regarding the cost of pre-training mutants while maintaining the computations of our experiment within a reasonable level. Indeed, the LeNet5 network architecture consists of seven layers containing basic deep learning modules including convolutional, average pooling, and fully connected layers. Consequently, the retraining process for generating each mutant is expected to be less time-consuming compared to deeper and more versatile network architectures such as ResNet20, which includes 20 layers. Therefore, performing the same experiment on subjects with larger DNNs and input sets is anticipated to require even more computational resources.
As described in Section II-C, the cost of mutation analysis is mainly due to performing two processes: mutant generation and MS calculation. TEASMA includes these two processes in its first step as described in Section III. Therefore, we repeated TEASMA’s first step utilizing DeepCrime’s pre-training MOs in place of DeepMutation++’s post-training MOs. Consequently, we can compare the cost and achieved Spearman correlation between FDR and MS based on mutants generated by DeepCrime and DeepMutation++ on identical randomly sampled input subsets of subject S1.
As described in Section II-B, when using pre-training MOs, different methods can be used for determining killed mutant and calculating MS. To conduct a comprehensive analysis we consider all the MS calculation variants for pre-training mutants detailed in Section II-B. As a result, we executed three variations of our cost analysis experiment, each differing in the way they calculate MS: (E4) $DeepCrimeMS$, (E5) $isKilledBasedMS$, and (E6) $DeepMutationMS$. Recall that E4 and E5 determine killed mutants based on Equation (5) while E6 does so based on Equation (9).
The duration of conducting mutation analysis using DeepMutation++ and DeepCrime on subject S1 is reported in Table VI. We must highlight that we adhere to the default procedure of each tool for filtering mutants with undesirable characteristics (equivalent, trivial, or redundant). The time taken for this procedure is included in the overall time of calculating MS. DeepCrime created 195 mutated versions of the training set or training program, retraining each version $20$ times (the tool’s default value), resulting in a total of 3900 mutant instances. As a result, Generating mutants using DeepCrime takes 43.52 hours. In contrast, DeepMutation++ directly mutated the original model and generated 304 mutated models in less than 10 seconds. Such a significant difference primarily results from the retraining process, and to some extent, is expected from any pre-training mutation testing tool. The overall time required for MS calculation based on pre-training mutants depends on the MS calculation approach and is 79.98 hours for E4 or E5 and 10.03 hours for E6. Both E4 and E5 use the $isKilled$ formula (Equation (5)) to determine killed mutants and calculate MS. In contrast, E6 follows the same MS calculation approach to DeepMutation++ (Equation (2)). Notably, the calculation of MS for the same set of input subsets, based on post-training mutants generated by DeepMutation++ and its way of calculating MS (E1) only takes 48.19 minutes.
TABLE VI: The total time of performing mutation analysis using DeepMutation++ and DeepCrime on subject S1
Tool DeepMutation++ DeepCrime
MS Formula E1 E4 or E5 E6
Mutant generation 9.10 s 43.52 h 43.52 h
MS calculation 48.19 m 79.98 h 10.03 h
Total 48.34 m 123.50 h 53.55 h
The time is reported in second (s), minutes (m), or hours (h)
The total time of performing mutation analysis based on the pre-training MOs of DeepCrime takes up to 123 hours when using DeepCrime’s statistical MS calculation approach (E4). This is approximately 153 times longer than the duration based on post-training MOs with DeepMutation++ and its original MS calculation (E1). Given the considerable time investment and the substantial computational resources needed for analyzing even a relatively small subject, the practicality of using pre-training MOs is limited in numerous application scenarios. Indeed, this total time is computed for the LeNet5 network, which is comparatively small compared to typical contemporary networks in industry. Utilizing a larger DNN network with more layers would increase mutant generation time due to the extended retraining time for each mutant. Similarly, larger input sets would contribute to an increased training time for each mutant. Consequently, as DNNs and input sets grow in size, the difference in the time required for performing mutation analysis based on pre- and post-training MOs is expected to increase.
To investigate whether the enormous cost of pre-training mutants would somehow be justified by improved FDR predictions, we also computed the Spearman correlation between FDR and MS based on E4, E5, and E6. This resulted in correlation values of 0.76, 0.61, and 0.98, respectively. We note that E4, representing the statistical approach proposed by DeepCrime for calculating MS, yields a much weaker correlation than E1 (0.74 vs. 0.99), the original way of calculating MS based on post-training mutants in DeepMutation++, as visible in Table III. Further, E6, which relies on the same formula as E1 but with pre-training MOs, shows a similar correlation to E1 with 0.98. Based on the above, we can conclude that there is no advantage, in terms of FDR prediction, to using pre-training MOs in DeepCrime. However, the execution of E6 required approximately 66 times the computation time of E1. Consequently, post-training MOs in DeepMutation++ clearly require much less computational resources for conducting mutation analysis. While this experiment, due to its high computational requirements, is conducted on a single subject of modest size, the findings underscore the substantial difference in computational resource requirements for mutation analysis based on pre- and post-training MOs. This observation, in addition to the strong association observed between mutants generated by post-training MOs and faults, further justifies the choice made in TEASMA to rely on post-training MOs.
VIII Conclusion

### v3

@@ V Results
Figure 5: Linear regression line between $\widehat{FDR}$ and $ActualFDR$ using MS
V-B RQ2: Which adequacy metrics lead to better FDR predictions on test sets?
To address this question, we analyze the relationship of the predicted FDR ($\widehat{FDR}$) with the actual one ($ActualFDR$) in the test set. To do so, for each subject, we selected a large number of random test subsets and measured the fit ($R^{2}$) and slope of a linear regression model between $\widehat{FDR}$ and $ActualFDR$. If the former can be trusted as an indicator of the latter, we should have a high $R^{2}$ and a slope close to 1. We should note that we repeat this process with each adequacy metric, consistently using the same set of test subsets. Table IV reports the results using all investigated test adequacy metrics. We can see a minimum $R^{2}$ of 0.94 and a slope that is either slightly below or above 1 when using MS, across all subjects. The results for DSC and IDC indicate a minimum $R^{2}$ value of 0.90. However, slopes tend to deviate more from 1 when compared to those observed with MS. Notably, for IDC, the slope is consistently greater than 1, suggesting that the prediction models built using this metric tend to underpredict FDR. LSC shows the lowest performance among investigated metrics, with a minimum $R^{2}$ value of 0.79 and slopes that consistently and significantly exceed 1. In particular, LSC shows substantial deviations, with slopes exceeding 2 for subject S2, 3 for subject S1, and 8 for subjects S4 and S6, indicating significant inaccuracies in FDR prediction for these cases.
For each metric, we calculate the average absolute deviation of the slopes of the linear regression lines between $\widehat{FDR}$ and $ActualFDR$ from 1 across all subjects. MS resulted in the lowest average deviation of 0.16, indicating that the slopes of the linear regression lines obtained using this metric across all subjects are closest to 1, on average. DSC has an average deviation of 0.34, significantly higher than MS, indicating it is less accurate than MS. IDC, with an average deviation of 0.44, is slightly higher than DSC but still relatively close. In contrast, LSC shows the largest average deviation of 2.56, indicating that its linear regression lines deviate the most from 1.
In summary, results indicate that the regression models developed on the training set using all investigated metrics, except for LSC, yield FDR predictions that are closely aligned with the actual FDR of the test sets. Therefore, FDR predictions obtained with TEASMA using MS, DSC, and IDC metrics provide a good basis on which to assess the adequacy of test sets and whether good accuracy results can be trusted to decide about deploying a DNN model.
The linear regression lines between $\widehat{FDR}$ derived from MS and $ActualFDR$, where the intercept is set to zero, are shown for all subjects in Figure 5. These regression lines clearly show a strong alignment between the two. The slopes of the lines fall within the range [0.95 1.23], thus indicating no strong bias in the FDR predictions. For most subjects, the prediction models tend to slightly underestimate the actual FDR, with the exception of subject S5, where a slight overestimation is observed. Due to space constraints, the linear regression lines for the other adequacy metrics are included in our replication package [44]. However, we report and discuss our observations here. DSC exhibits the widest range of slope variations, from 0.81 for subject S5 to 1.67 for subject S3. While the prediction models tend to underestimate the actual FDR for most subjects, they overestimate it for three subjects. In contrast, IDC is the only metric where the regression slopes consistently exceed 1, ranging from 1.10 to 1.62. This suggests that for test subsets, detecting all faults is easier than achieving maximum IDC coverage.
Answer to RQ2: A linear regression model between $\widehat{FDR}$ and $ActualFDR$ on the test set can be fitted with a minimum $R^{2}$ value of 0.94 using MS, and 0.90 using DSC, and IDC, with slopes slightly below or above 1. This indicates that, for each subject, the selected regression model built on the training set using these metrics, accurately predicts the FDR of a test set based on its AS. Consequently, TEASMA provides a reliable methodology for assessing test sets and whether to trust high accuracy results when validating a DNN model.
TABLE V: Accuracy of the FDR predictions using accurate adequacy metrics
Accuracy Adequacy Subjects Avg
metric metric S1 S2 S3 S4 S5 S6 S7 S8
RMSE MS (E1) 0.12 0.11 0.12 0.09 0.05 0.07 0.10 0.09 0.09
DSC 0.21 0.24 0.26 0.06 0.16 0.16 0.08 0.10 0.17
IDC 0.20 0.23 0.25 0.15 0.06 0.23 0.20 0.15 0.18
V-C RQ3: How accurately can we predict a test set’s FDR?
To address this question, we compare the adequacy metrics identified in the previous question as accurate predictors of FDR: MS, DSC, and IDC. To evaluate the prediction accuracy of each metric, we employ RMSE, a standard and easy to interpret measure for assessing prediction accuracy. RMSE measures the average magnitude of the errors between predicted and actual values. It computes the square root of the average squared differences between predicted and actual values, making it sensitive to larger errors and penalizing them more than smaller ones. This characteristic of RMSE is particularly valuable in our context, where large errors between the predicted FDR and the actual FDR are especially undesirable. Table V presents the RMSE values, between $\widehat{FDR}$ and $ActualFDR$, for each metric. MS achieves the lowest RMSE among all metrics for most subjects (five out of seven) and has the second-lowest RMSE, with only a small difference from the lowest value, for the remaining two subjects, indicating that MS is consistently accurate across different subjects. The average RMSE for MS is 0.09, indicating that, on average, the difference between the predicted FDR and the actual FDR is 0.09. This low value suggests that predictions using MS are quite accurate. In contrast, the average RMSE values for DSC and IDC are 0.17 and 0.18, respectively, nearly double that of MS. These higher RMSE values imply that DSC and IDC are significantly less accurate at predicting FDR compared to MS.
While the shape of the relationship between IDC and FDR is linear for some subjects and close to linearity for others, indicating it could be an accurate metric, these results show that IDC is not the most accurate for predicting FDR. Our investigation reveals that the shape of the relationship between IDC and FDR for test subsets is not similar to that of training subsets. One explanation for this can be that, unlike other metrics, calculating IDC is entirely independent of the DNN model and is solely based on input features. Consequently, inputs that are correctly predicted by the DNN model–and thus do not contribute to fault detection–can still increase the IDC coverage. Given that the accuracy of the DNN model on test sets is typically lower than on its training set, the proportion of correctly predicted inputs that can increase coverage without contributing to FDR is higher in the training subsets than in the test subsets. This discrepancy affects the shape of the relationship between IDC and FDR on the training subsets, making it different from the relationship on the test subsets. It is important to note that the relationship between IDC and FDR on the test subsets is mostly non-linear. For subjects where a linear relationship exists for both training and test subsets, the slope of the linear regression line is higher for the test subsets. This indicates that, for training and test subsets with identical IDC coverage, the test subsets consistently achieve higher FDR.
Answer to RQ3: Its relatively low RMSE (9% on average) across all subjects, indicates that MS calculated based on DeepMutation’s post-training MOs is a good solution for accurately predicting FDR, following the TEASMA methodology, and fares better than solutions relying on DSC and IDC.
V-D Fault clusters
In this section, we report on the fault clusters we used in our experiments since they play a central role in our experiments. These clusters are obtained by applying the approach described in Section II-D on the DNN model’s training set as outlined in Section III-B. This clustering approach, like many others, relies on several hyperparameters such as the number of neighbors considered locally connected in UMAP and the minimum size of clusters in HDBSCAN. These hyperparameters should be effectively fine-tuned to achieve the best clustering results. Therefore, following the original implementation [18], we applied this approach to each subject using several hyperparameter configurations. We then selected the optimal configuration through the same manual and metric-based evaluations introduced in the original paper [18].
As described in Section II-D, the clusters obtained for each subject were evaluated based on the Silhouette score. For each subject, the configuration that achieved the highest Silhouette score was selected to obtain the final clusters. Table VI provides detailed information on the fault clusters obtained for each subject in our experiments, including the total number of fault clusters as well as their Silhouette score and DBCV value. Achieving a Silhouette score close to 1 is rarely feasible in practice but, with a minimum Silhouette score of 0.60 across all subjects, we seem to achieve good clustering based on what is reported in the literature [31, 18]. This score suggests that the clusters are well-defined and distinct, reflecting a good balance between intra-cluster cohesion and inter-cluster separation. The clusters’ DBCV scores, which exceed 0.5 for most subjects, provide additional evidence supporting the quality of the clustering results. These scores further suggest that the final clusters are not only well-separated but also dense, which is indicative of high-quality clustering.
TABLE VI: Information about fault clusters
Subjects S1 S2 S3 S4 S5 S6 S7 S8
#Faults 307 354 191 512 607 590 718 3118
Silh. 0.88 0.70 0.69 0.79 0.66 0.78 0.78 0.60
DBCV 0.81 0.51 0.52 0.66 0.51 0.66 0.66 0.47
(a) $A\>final\>cluster$
(b) $A\>noisy\>cluster$
Figure 6: Heatmap examples of the fault clusters
Furthermore, we conducted an exhaustive manual evaluation of the features for each cluster as described in the original paper [18]. For this purpose, we evaluate features’ heatmaps related to final clusters. Two representative examples of these heatmaps are depicted in Figure 6. Each heatmap’s rows correspond to the input IDs within a cluster, while columns represent their respective features, with colors indicating feature values. The first heatmap, in Figure 6(a), is related to a quality cluster characterized by well-clustered inputs sharing common feature distribution patterns. Conversely, the second heatmap, in Figure 6(b), illustrates a cluster containing poorly-clustered inputs. We observed that most of the clusters look like Figure 6(a), demonstrating coherent feature distribution patterns and indicating effective clustering.
V-E Threats to Validity
@@
Internal threats to validity. One of the internal threats to validity pertains to our sampling approach, as our results are based on sampled input subsets. To address this, we selected a large number of randomly sampled input subsets of varying sizes. Furthermore, when experimenting with MS, we employed uniform sampling in addition to random sampling, since one of the MS calculation formulas used in our experiments considers the number of images from each input class in a subset. Another internal threat to validity stems from selected parameters and configurations used throughout the experiments. This encompasses parameters for clustering mispredicted inputs to estimate faults and for configuring the calculation of AS. To mitigate this concern, regarding clustering parameters, we explored different settings for each subject and opted for the best configuration based on the evaluation of the Silhouette index [31] and DBCV scores [32], as suggested by Aghababaeyan et al. [18]. For MO configuration, we used the default settings of the DeepMutation++ tool, which yielded accurate results in terms of FDR prediction. For calculating DSC and LSC, following the original study [6], we used the deepest layer of the model and set the number of buckets to 1000. The lower and upper bounds were set using the SA values observed during training. For calculating IDC, we used the publicly available VAEs for subjects sharing identical input sets with those used in the original study [5] and trained new VAEs only for the remaining input sets.
@@
Construct threats to validity. A potential threat is related to the calculation of MS, SC, and FDR. To mitigate this concern and ensure the robustness of our results, we considered multiple ways to compute MS, as described in Section II-A3. Regarding the SC calculation, it is important to note the implementations we employed. We primarily utilized the original implementations of DSC and LSC provided by Kim et al. [6]. However, due to the substantial computational cost associated with calculating LSC on our largest input set, ImageNet, we opted for an alternative implementation provided by Weiss et al. [49] for this specific dataset. It has been thoroughly evaluated in the corresponding study by Weiss et al. and has shown comparable outcomes to the original implementation. We thus ensure the feasibility of our experiments while maintaining the integrity of our results. For FDR, we leveraged a SOTA fault estimation approach [18] based on clustering mispredicted inputs. This approach has been thoroughly validated and the resulting clusters have been carefully analyzed by the authors. We reused their publicly available approach to achieve accurate fault identification.
@@
External threats to validity. To address potential concerns regarding the generalizability of TEASMA, we performed our experiment on multiple subjects including various combinations of seven DNN models with different internal architectures and six input sets containing diverse types of images. The selected input sets include large input sets, such as Cifar-100 and ImageNet, to assess the practicality and scalability of the proposed methodology. Our results show great consistency in terms of FDR prediction accuracy across all these models and input sets, further validating the generalizability of TEASMA. Another potential threat may arise from the specific adequacy metrics selected for evaluating TEASMA. To mitigate this threat, we investigated four SOTA metrics, including the widely-used DSC and LSC metrics and the recent input feature coverage IDC metric, as well as mutation analysis. Another threat might be related to the specific set of MOs that we considered in our experiment with MS. For practical reasons, we focused on post-training MOs and investigated all the MOs implemented by the DeepMutation++ tool, also considering whether subsets of MOs might yield better results. Other MOs may yield different results in the future.
VI Related Work
In this section, we review prior work related to our proposed methodology from two perspectives: test selection and generation, and test adequacy assessment, within the context of DNN models.
VI-A Test Selection and Generation
Testing and enhancing a DNN model requires large labeled input sets. In many contexts, however, the scarcity of labeled inputs poses a significant challenge, especially when these models are applied in and fined-tuned to different deployment contexts. Consequently, recent research has focused on developing innovative methods for selecting unlabeled test inputs to not only test (after labeling) but also improve the model via retraining with the selected set [50, 51].
Several selection methods assign a score to each test input that captures its priority for getting labeled and tested, enabling testers to select test inputs with higher scores based on their labeling budget. Unlike surprise adequacy metrics introduced by Kim et al. [7], such as DSA, LSA, and MDSA, which measure how surprising or unexpected a new test input is to a DNN model given its training input set (as detailed in Section II), uncertainty-based metrics, such as DeepGini [27], Margin [52], and MaxP [52], estimate the model’s confidence in predicting each test input. They aim to prioritize inputs on which the model is least confident in its predictions and that are thus most informative or challenging for a model. Similarly, entropy-based methods [53] compute the entropy of the predicted class probabilities for each test input and prioritize inputs with higher entropy, indicating greater uncertainty in the model’s predictions.
Other test selection approaches, rather than scoring individual inputs, select a test subset with a predefined labeling budget. For instance, Aghababaeyan et al. [54] proposed DeepGD, a black-box multi-objective search-based approach. DeepGD selects a fixed-size subset of inputs guided by the uncertainty of model output probabilities and the diversity of selected inputs’ features. It has been shown to be able to select test sets capable of effectively revealing diverse faults in the DNN model and guiding the enhancement of the model through retraining with the selected inputs. Ma et al. [51] conducted an empirical study with various selection metrics (including DSA, DeepGini, and the entropy-based metric) and random selection to investigate the impact of various ranges of distribution shift on their effectiveness for model enhancement. Their observations revealed that no individual metric consistently outperforms the others across all distribution shift ranges. Building on these insights, they introduced a new approach for selecting a fixed-size subset of inputs, which mitigates the effects of distribution shifts and exhibits up to a 30% improvement in model accuracy. Evaluating the performance of test selection methods on test sets with a distribution shift from the training set is particularly important for methods that aim to enhance the model in new deployment contexts through retraining. In this paper, however, our focus is on assessing the adequacy of test sets. In scenarios where a distribution shift occurs between the training set and the deployment environment and leads to poor model accuracy, model retraining is necessary and test adequacy assessment is not useful as one cannot have confidence in the model regardless of the test set used. After proper retraining, test adequacy assessment can be performed on the retrained model to validate its high-accuracy test results. In our experiments, to prevent any potential biases, we intentionally relied on the subject models’ original test sets since we had no involvement in their selection.
Zohdinasab et al. [55] introduced DeepHyperion-CS, a tool for generating inputs that explores the feature space to cover unseen feature combinations or improve performance compared to similar existing inputs. However, the study did not investigate the potential of using feature space coverage as a metric for test adequacy or its correlation with a test set’s fault detection capability. Selecting appropriate features is crucial when employing feature-based coverage metrics, as their effectiveness hinges on choosing features that accurately represent significant system properties. Zohdinasab et al. [55] aim to provide an interpretable characterization of DL systems by identifying which features of the test inputs influence the system’s incorrect behavior. Nonetheless, identifying such features is challenging and involves a systematic approach that requires domain experts’ assistance. This manual process is time-consuming and labor-intensive, requiring human assessors to analyze existing inputs and select relevant features for a given domain. In contrast, Attaoui et al. [34] proposed a feature-based test selection approach for the image domain that relies on a pre-trained model to automatically extract features of the error-inducing inputs. Consequently, they can detect the root causes of mispredictions by clustering these inputs based on their features. Finally, they use these clusters to effectively retrain and enhance the performance of the DNN model.
Regardless of the selection or generation approach, assessing the adequacy of a selected test set is essential to mitigate the risk of insufficient testing. Based on our findings, TEASMA can accurately assess the adequacy of a test set and thus confidently verify whether high-accuracy test outcomes can be trusted.
VI-B Test Adequacy Assessment
Test adequacy assessment refers to the process of evaluating the quality of a test set and its capacity to validate test outcomes [56]. In traditional software testing, coverage metrics [57] and mutation testing [56] stand out as pivotal methods for assessing test adequacy. Coverage metrics, including statement, branch, condition, and path coverage, measure the comprehensiveness of a test suite and its ability to execute different parts of the system code. Similarly, researchers have proposed various neuron coverage metrics in the context of DNN models for test assessment [8, 9, 10, 11, 20]. They aim to measure how well the test set covers the model’s internal neuron activations, ensuring that a wide range of the DNN model’s behaviors are examined during testing. However, these metrics have shown no significant correlation with the number of mispredicted inputs or detected faults in a test set [19, 58]. Consequently, in our experiments to evaluate TEASMA’s methodology with coverage metrics, we relied on SOTA surprise coverage metrics, DSC and LSC, alongside the recently introduced black-box coverage metric IDC, as described in Section II.
Mutation testing has been established as a widely-used technique for assessing test suites in traditional software [56, 16]. As described in detail in Section II, this technique has been adapted for DNNs and specific MOs have been proposed. The potential of these operators to address the high cost of labeling test inputs in DNNs by guiding test input generation and prioritization has been investigated in the literature [59, 60]. Riccio et al. [59] developed DeepMetis, a search-based input generation tool that utilizes the mutants generated by DeepCrime’s pre-training operators to guide the process of test input generation. Their goal is to generate inputs that can augment an existing test set and increase its capability to kill pre-training mutants. Given that DeepCrime’s operators are based on real faults [14], test sets that kill more mutants are assumed to be able to detect more faults. Wang et al. [60] proposed PRIMA, a test input prioritization approach based on mutation analysis. In addition to four post-training MOs similar to DeepMutation’s operators, they define input MOs for some specific domains including text and image inputs. PRIMA prioritizes inputs that are able to kill many mutated models based on the assumption that such inputs are more likely to reveal faults. Additionally, it prioritizes inputs that produce different prediction results with many mutated versions. The rationale is that if a slight mutation on a test input alters the prediction results, it indicates that the mutated part is actively utilized by the model, making it sensitive to capturing errors in the model. Both DeepMetis [59] and PRIMA [60] aim to achieve test sets with higher MS. However, neither of these approaches investigate the correlation between MS and FDR, nor do they intend to help assess the adequacy of selected or generated test sets.
The application of post-training operators for adversarial sample detection in DNNs has also been investigated by Wang et al. [61]. They introduced a new approach relying on DeepMutation’s MOs, based on the observation that adversarial samples are more sensitive to mutations in the DNN model compared to normal samples. This sensitivity is leveraged to distinguish between normal and adversarial samples at runtime.
Despite the availability of various test adequacy metrics such as MS, DSC, LSC, and IDC, a critical gap remains in the development and evaluation of a comprehensive methodology for test adequacy assessment. TEASMA addresses this gap by providing a reliable and validated methodology to evaluate test set adequacy, regardless of the selection or generation technique employed, utilizing and comparing SOTA test adequacy metrics. Our results show that TEASMA when deployed with DSC, IDC, and MS calculated based on post-training MOs, provides an accurate prediction of the fault detection ability of test sets and therefore their adequacy. MS seems to yield the most accurate results.
VII Conclusion